A software designer is mapping the streets for a new racing game. All of the streets are depicted as either perpendicular or parallel lines. The equation of the lane passing through A and B is -7x + 3y = -21.5. What is the equation of the central street PQ?


1
Expert's answer
2021-07-15T06:14:28-0400

Solution.



The equation of the line passing through A and B is: "-7x + 3y = -21.5."

It can be written as: "y=\\frac{7}{3}x\u200b\u2212\\frac{21.5}{3}.\n\u200b"

By comparing with the standard form "y=mx+c," the slope of this line is "m=\frac{7}{3}."

Central street PQ is perpendicular to the line passing through A and B. Product of slopes of perpendicular lines is "m1*m2=-1".

Slope of the perpendicular line is:"Slope\; of \;perpendicular\; line=\newline = \frac{-1}{slope\;of\;parallel\;line}=\frac{-1}{\frac{7}{3}}=\frac{-3}{\;7}."

Thus the equation of line is: "y=mx+c; \newline
y=\frac{-3}{7}x+c; \newline
7y= -3x+7c; \newline
7y+3x=7c."

Dividing by 2:

"3.5y+1.5x=3.5c."

To find c with just this information is impossible.It is needed to check the options of the line with the slope -3/7

on the figure.

The line PQ passes through the point (7,6) on the figure.Hence the equation of the line can be found out, using point slope form of a line: "y\u2212y_1=m(x\u2212x_1)."

Slope is "\frac{-3}{7}" .

"(x_1,y_1) \;is\; (7,6)."

"y-6=\frac{-3}{7}(x-7); \newline
7(y-6)=-3(x-7); \newline
7y-42=-3x+21; \newline
7y+3x=21+42; \newline
7y+3x=63."

This is the equation of the central line.

Dividing by 2:

"3.5y+1.5x=31.5."

Or

"-3.5y-1.5x=-31.5" is also the equation.

Answer:

"3.5y+1.5x=31.5"

or

"-3.5y-1.5x=-31.5".